A regression between the sales (Y) and the monthly advertising expenditures (X) resulted in the following predicted regression equations:

Y with hat on top = 10.9 + 0.23 x

One observation in the sample data set shows that one store spend $68 K in advertising with sales of $29 K . What is the predicted sales for this store?

$ 33 K

$26.54 K

$17.57 K

$29.03 K
Answer #1

Here, it is given that

Y =Sales

X = Monthly advertising expenditures

The regression equation is û = 10.9 +0.23.

If one store spend $68K in advertising with sales of $29K. i.e 89 = 1

The predicted sales for this store is

y = 10.9+0.23 * 68 9 = 10.9+15.64 û = 26.54K

The predicted sales for this store is $26.54 K.